# DCM: Dual-Expert Consistency Model for Efficient and High-Quality Video Generation

### v1

@@ 2.1 Diffusion Models For Video Synthesis
Video Diffusion Models have witnessed rapid advancements with diffusion models [8, 9, 2]. Building on the Diffusion Transformer (DiT) [41] pre-training, a notable breakthrough is the development of high-fidelity video diffusion models [40, 19, 51, 38, 20, 10, 59, 18]. However, scaling these models for long videos incurs significant training and inference costs. Recently, LTX-Video [7] designed Video-VAE that achieves a high compression ratio for efficient self-attention. Pyramid flow [13] introduced a unified pyramidal flow matching algorithm for efficient video generative modeling. Moreover, while efficient fast diffusion samplers [48, 25, 29, 30, 15] reduce inference steps, further reduction often severely degrades performance. Diffusion distillation offers a promising way to further minimize sampling steps while maintaining the visual quality.
@@ 2.2 Diffusion Model Distillation
Trajectory-preserving distillation methods exploit the fact that diffusion models learn an Ordinary Differential Equation (ODE) trajectory and aim to predict the exact teacher output in fewer steps. Among the earliest studies on diffusion distillation, Luhman et al. [31] and DSNO [64] proposed training the student model using noise-image pairs precomputed by the teacher model with an ODE solver. Progressive distillation [45, 39] reduces the final number of sampling steps by iteratively applying the distillation process to halve the number of sampling steps of previous model. Instaflow [26, 27] progressively learns straighter flows, enabling accurate one-step predictions over larger distances. Consistency models [50, 49, 32, 33, 24, 65, 28], BOOT [5] and TRACT [1] learn to map samples along the ODE trajectory to another point to achieve self-consistency. The consistency trajectory model [16] was designed to mitigate discretization inaccuracies and accumulated estimation errors in the multistep consistency model sampling. PCMs [53] phase the ODE trajectory into several sub-trajectories and only enforce the self-consistency property on each sub-trajectory, thus alleviating the limitations of CMs. Trajectory-preserving distillation enables stable optimization but can degrade visual quality, leading to blurriness or distortions when sampling with fewer steps.
@@ 3.1 Preliminary
$\mathcal{L}_{DM}=\mathbb{E}_{\boldsymbol{x},\boldsymbol{\epsilon}\sim\mathcal{N}(0,1),t}\big{[}||\boldsymbol{\epsilon}-\boldsymbol{\epsilon}_{\theta}(\boldsymbol{x}_{t},t)||^{2}_{2}\big{]}.$ (2)
@@ 3.4 Expert-specific Optimization Objective
$\mathcal{L}_{TC}$ $=||(\boldsymbol{x}^{t_{\kappa}}_{l:L}-\boldsymbol{x}^{t_{\kappa}}_{0:L-l})-(\hat{\boldsymbol{x}}^{t_{\kappa}}_{l:L}-\hat{\boldsymbol{x}}^{t_{\kappa}}_{0:L-l})||^{2}_{2}.$
@@ 4.1 Experimental Setup
Implementation Details For HunyuanVideo, we selected trajectories with 50 Euler steps and used the default sampling parameters from diffusers. The distillation was conducted on 129-frame video sequences at a resolution of $1280\times 720$ with a batch size of 6. For the semantic expert denoiser, we performed $1000$ iterations of distillation with a learning rate of $61e-$ 6, while for the details expert denoiser, we trained for $1000$ iterations with a learning rate of $65e-$ 6. For CogVideoX-2B, we selected trajectories with 50 DDIM steps. The distillation was conducted on 29-frame video sequences at a resolution of $720\times 480$ with a batch size of 4. In the first-stage fine-tuning, we distilled for approximately 1000 steps, while in the second-stage fine-tuning, we distilled for around 500 steps, both with a learning rate of $61e-$ 6. All experiments were conducted on 24 NVIDIA A100 80GB GPUs.
@@ 4.3 Ablation Study
Effect of Optimization Decoupling (OD) Through Experiments (1) and (2), we observe that decoupling the optimization of semantic and detail modeling significantly improves the semantic and quality scores of the synthesized videos. As shown in Fig. 7, the optimized decoupled model synthesizes videos with better semantic and detail quality, where the motion of characters and facial details appear more natural.

### v2

# Dual-Expert Consistency Model for Efficient and High-Quality Video Generation

@@ 2.1 Diffusion Models For Video Synthesis
Video Diffusion Models have witnessed rapid advancements with diffusion models [8, 9, 2], driving a wide range of visual content generation applications[70, 12, 71, 13, 72]. Building on the Diffusion Transformer (DiT) [45] pre-training, a notable breakthrough is the development of high-fidelity video diffusion models [44, 21, 55, 42, 22, 10, 63, 20]. However, scaling these models for long videos incurs significant training and inference costs. LTX-Video [7] designed Video-VAE that achieves a high compression ratio for efficient self-attention. Pyramid flow [15] introduced a unified pyramidal flow matching algorithm for efficient video generative modeling. The acceleration method based on sparse attention [68, 69] and feature cache [73, 41] has also improved inference efficiency. Moreover, while efficient fast diffusion samplers [52, 28, 32, 33, 17] reduce inference steps, further reduction often severely degrades performance. Diffusion distillation offers a promising way to reduce sampling steps while maintaining visual quality.
@@ 2.2 Diffusion Model Distillation
Trajectory-preserving distillation methods exploit the fact that diffusion models learn an Ordinary Differential Equation (ODE) trajectory and aim to predict the exact teacher output in fewer steps. Among the earliest studies on diffusion distillation, Luhman et al. [34] and DSNO [74] proposed training the student model using noise-image pairs precomputed by the teacher model with an ODE solver. Progressive distillation [49, 43] reduces the final number of sampling steps by iteratively applying the distillation process to halve the number of sampling steps of previous model. Instaflow [29, 30] progressively learns straighter flows, enabling accurate one-step predictions over larger distances. Consistency models [54, 53, 35, 36, 27, 75, 31], BOOT [5] and TRACT [1] learn to map samples along the ODE trajectory to another point to achieve self-consistency. The consistency trajectory model [18] was designed to mitigate discretization inaccuracies and accumulated estimation errors in the multistep consistency model sampling. PCMs [57] phase the ODE trajectory into several sub-trajectories and only enforce the self-consistency property on each sub-trajectory, thus alleviating the limitations of CMs. TCM [23] generalizes consistency training to the truncated time range to prevents the truncated-time training from collapsing to a trivial solution. Trajectory-preserving distillation enables stable optimization but can degrade visual quality, leading to blurriness or distortions when sampling with fewer steps.
@@ 3.1 Preliminary
[…]
@@ 3.4 Expert-specific Optimization Objective
[…]
@@ 4.1 Experimental Setup
Implementation Details For HunyuanVideo, we selected trajectories with 50 Euler steps and used the default sampling parameters from diffusers. The distillation was conducted on 129-frame video sequences at a resolution of $1280\times 720$ with a batch size of 6. For the semantic expert denoiser, we performed $1000$ iterations of distillation with a learning rate of 1e61e-6, while for the details expert denoiser, we trained for $1000$ iterations with a learning rate of 5e65e-6. For CogVideoX-2B, we selected trajectories with 50 DDIM steps. The distillation was conducted on 29-frame video sequences at a resolution of $720\times 480$ with a batch size of 4. In the first-stage fine-tuning, we distilled for approximately 1000 steps, while in the second-stage fine-tuning, we distilled for around 500 steps, both with a learning rate of 1e61e-6. All experiments were conducted on 24 NVIDIA A100 80GB GPUs.
@@ 4.3 Ablation Study
Effect of Optimization Decoupling (OD) Through Experiments (1) and (2), decoupling the optimization of semantic and detail modeling significantly improves the semantic and quality scores. As shown in Fig. 7, the optimized decoupled model synthesizes videos with better semantic and detail quality, where the motion of characters and facial details appear more natural.
@@ 5 Conclusion and Discussion
6 Acknowledgement
This study is supported by the Ministry of Education, Singapore, under its MOE AcRF Tier 2 (MOE-T2EP20221-0012, MOE-T2EP20223-0002), and under the RIE2020 Industry Alignment Fund – Industry Collaboration Projects (IAF-ICP) Funding Initiative, as well as cash and in-kind contribution from the industry partner(s).